OpenStudy (anonymous):

A boat travels 12 km upstream and back in total time of 1 hr and 15 minutes. When there's constant current speed is 1.5 km/h, find speed of the boat in still water.

OpenStudy (baru):

\(t_1\): time taken for upstream travel \(t_2\):time taken for downstream travel speed of boat: s downstream speed: s+1.5 upstream speed s-1.5 \(t_1=12/(s-1.5)\\t_2=12/(s+1.5)\) \(t_1+t_2= 1hr15min\)

OpenStudy (anonymous):

you calculated how to find speeds in downstream and upstream, but what about speed in still water?

OpenStudy (baru):

i have not calculated anything...I have left the last step for you to solve. just substitute \(t_1~and~t_2\) into the last equation. solve for "s" s is the speed of boat in still water

OpenStudy (anonymous):

Right. Thanks a lot for the help! :)